# CoSPED: Consistent Soft Prompt Targeted Data Extraction and Defense

### v2

@@ Abstract
Large language models have gained widespread attention recently, but their potential security vulnerabilities, especially privacy leakage, are also becoming apparent. To test and evaluate for data extraction risks in LLM, we proposed CoSPED, short for Consistent Soft Prompt targeted data Extraction and Defense. We introduce several innovative components, including Dynamic Loss, Additive Loss, Common Loss, and Self Consistency Decoding Strategy, and tested to enhance the consistency of the soft prompt tuning process. Through extensive experimentation with various combinations, we achieved an extraction rate of 65.2% at a 50-token prefix comparison. Our comparisons of CoSPED with other reference works confirm our superior extraction rates. We evaluate CoSPED on more scenarios, achieving Pythia model extraction rate of 51.7% and introducing cross-model comparison. Finally, we explore defense through Rank-One Model Editing and achieve a reduction in the extraction rate to 1.6%, which proves that our analysis of extraction mechanisms can directly inform effective mitigation strategies against soft prompt-based attacks.
@@ Introduction
With the growing popularity of ChatGPT (Wu et al. 2023; Yan et al. 2023), Large Language Models (LLMs) (Devlin et al. 2019; Radford et al. 2019; Raffel et al. 2020; Chang et al. 2024) have become one of the most impactful developments in deep learning.
However, their potential security problems, especially privacy leakage, are attracting increasing attention (Yao et al. 2024; Carlini et al. 2021, 2022; Ozdayi et al. 2023; Zhang, Wen, and Huang 2023). Recent works indicate that LLMs tend to memorize training data, resulting in considerable potential for leakage (Carlini et al. 2021, 2022), which raises an urgent need for deeper analysis and effective mitigation strategies.
@@
Soft prompts (Lester, Al-Rfou, and Constant 2021) are trainable continuous vectors that can guide model behavior without altering the model architecture (Li and Liang 2021; Wu et al. 2024). While typically used for task adaptation, they can also be exploited for data extraction attacks.
Prior work has utilized soft prompts to conduct data extraction attacks in white-box settings (Zhang, Wen, and Huang 2023; Ozdayi et al. 2023). The Controllable Language Model (CLM) framework (Ozdayi et al. 2023) offers an initial treatment of this problem by training soft prompts for extraction and proposing a coarse-grained defense. The Ethicist system (Zhang, Wen, and Huang 2023) extends this line of work with a smoothing loss and calibrated confidence estimation to prioritize high-risk tokens.
Previous works (Ozdayi et al. 2023; Zhang, Wen, and Huang 2023) have explored soft prompt-based extraction. We observe that generation-based extraction methods often produce highly variable or incoherent outputs across repetitions, loss configurations, and decoding choices. However, they face three key limitations: (1) lack of consistent mechanisms for soft prompt tuning, (2) limited extraction ability, and (3) evaluation on single structural models only.
@@
  • We conduct extensive empirical studies on multiple open-source LLMs, including both the GPT-Neo series and the Pythia series, introducing model structural related influence.
  • We implement and evaluate a defense mechanism based on the ROME model editing, demonstrating its effectiveness in mitigating soft prompt-based extraction attacks.
@@
Large-scale pre-trained language models often unintentionally memorize training data, leading to privacy risks where sensitive information can be leaked (Carlini et al. 2021, 2022; Tirumala et al. 2022).
Various studies have shown that models can inadvertently store and reveal personal information, which can be extracted through targeted queries (Carlini et al. 2019, 2021, 2022; Lehman et al. 2021; Kim et al. 2023).
Defense efforts include data deduplication (Kandpal, Wallace, and Raffel 2022), sensitive information deletion (Patil, Hase, and Bansal 2023a), and differential privacy (Abadi et al. 2016; McMahan et al. 2017), but these strategies remain insufficient against evolving attacks (Yao et al. 2024).
Extraction attacks focus on extracting LLM’s memorized data, including model stealing (Kariyappa, Prakash, and Qureshi 2021; Truong et al. 2021), gradient leakage (Li et al. 2023), and training data extraction (Carlini et al. 2021; Zhang, Wen, and Huang 2023; Ozdayi et al. 2023; Yang et al. 2024). This work focuses on the latter, which poses critical risks by extracting sensitive training data (Kim et al. 2023; Yang et al. 2024; Zhang, Wen, and Huang 2023; Ozdayi et al. 2023).
@@
Prompt tuning (Lester, Al-Rfou, and Constant 2021) optimizes continuous input embeddings (soft prompts) to guide model behavior without modifying model parameters (Li and Liang 2021). Unlike text prompts, soft prompts directly manipulate input at the embedding level (Lester, Al-Rfou, and Constant 2021).
Recent works have exploited soft prompts for data extraction attacks. CLM (Ozdayi et al. 2023) trains soft prompts to extract memorized suffixes given prefixes, achieving improved extraction rates over baselines. However, CLM’s defense method lacks robustness and its training improvements are limited.
Ethicist (Zhang, Wen, and Huang 2023) enhances extraction with a smoothing loss that optimizes generation probabilities for high-loss tokens and uses calibrated confidence estimation for suffix selection. However, Ethicist only considers top-k highest loss tokens, ignoring others, and evaluates solely on GPT-Neo 1.3B.
To address these limitations, we propose enhancements including: (1) multiple loss components capturing all token loss trends, (2) novel decoding strategies for consistent outputs, and (3) evaluation across multiple model families with transferable datasets. Additionally, we explore defense mechanisms with ROME (Meng et al. 2023a) model editing and how understanding these extraction patterns enables more targeted and effective defenses.
@@
[…]
We recognize LLM’s training dataset as $D=\{X_{i}(P_{i},S_{i})\}$, which includes sets of $X_{i}$ sequences divided as pairs of prefixes $P_{i}$ and suffixes $S_{i}$. We define the attacker-side training split as
$D_{a}=\{X_{i}(P_{i},S_{i})\mid X_{i},P_{i},S_{i}\in D,1\leq i\leq|D|\}.$
$D_{a}$ enumerates every prefix–suffix pair available for soft prompt tuning, ensuring that optimization covers the same distribution as the memorized corpus. $P_{i}$ and $S_{i}$ consist of $X_{i}$ from the prepared dataset $D$ and $i$ is ranging inside $|D|$. This is a reasonable assumption given that large-scale pre-trained data often includes publicly available datasets such as Books3 (Zhu et al. 2015), Common Crawl, and The Pile (Gao et al. 2020). $D_{a}$ is utilized to train the attacking model to tune the soft prompt, and the goal is to predict suffixes for generations.
It can be certain that testing dataset $D_{p}$ will not overlap with $D_{a}$, although $D_{a}$ and $D_{p}$ both come from LLM’s training data $D$. The main reason is their different training and testing characteristics.
@@
To perform data extraction with a soft prompt, we initialize soft prompt $\mathcal{Z}$ with $K\times d$ vector parameters, where $K$ is the soft prompt length and $d$ is the model’s embedding dimension. During training, $\mathcal{Z}$ is concatenated with prefixes $P_{i}$ from $D_{a}$ and optimized to minimize the loss $\mathcal{L}$ for generating corresponding suffixes $S_{i}$, while keeping the LLM parameters frozen. During evaluation, the trained soft prompt is used with prefixes from $D_{p}$ to generate suffixes $S_{p}$, which are compared with ground truth using exact match metrics. Figure 2 illustrates this process.
@@
We study a targeted training-data extraction adversary with full white-box access to the victim language model. The attacker can inspect model parameters, gradients, and hidden states, and can load the model on hardware they control. Their goal is to reconstruct memorized suffixes $S$ from the training corpus given a prefix $P$ that is known or guessed. Consistent with prior literature, we assume the attacker can fine-tune continuous soft prompts $\mathcal{Z}$ but does not modify the model weights. The adversary may repeatedly query the model during training and decoding, select loss functions, and post-process generations, but they cannot alter the underlying training dataset or access user-specific runtime logs. Our defense analysis considers the same threat actor and evaluates how editing internal representations can reduce the attack success rate without degrading benign utility.
@@
LLM, in basic, provides a maximum likelihood estimation (MLE) loss.
Other than MLE, our method explored several losses, including smooth loss, focal loss, dynamic loss, additive loss, and common loss, some of which consist of the best $\mathcal{L}_{\text{extra}}$ shown in Fig. 2, which can either enhance or replace the original MLE loss.
$S$ represents the suffix sequence the model is trying to generate, and the MLE loss is calculated based on each $t_{i}$ token loss, which is influenced by the soft prompts $\mathcal{Z}$, input prefix $P$, and previously generated tokens $t_{<i}$.
(Zhang, Wen, and Huang 2023) proposed an extra loss function called Smooth loss, which is meant to increase the top $N$ lowest loss values and additionally optimize the generation probabilities for these $N$ tokens. It gives improvements while combining with original $\mathcal{L}_{\text{MLE}}$, but in our method, we are exploring more combinations of different losses, including smooth loss, to test for more potential.
@@
Focal loss(Lin et al. 2018) as presented in Eq. 1 is a modification of the standard cross-entropy loss, primarily designed to address the class imbalance problem. In our scenario, focal loss focuses on penalizing hard-to-classify tokens more than easier tokens, thereby emphasizing learning from difficult instances.
$\mathcal{L}_{\text{Focal}}=-\frac{1}{|S|}\sum_{i=1}^{|S|}\alpha_{t}(1-p_{i})^{\gamma}\log p_{i}.$ (1)
Equation 1 reweights each token-level log likelihood with $\alpha_{t}$ and the focusing parameter $\gamma$ to emphasize hard examples. The probability term satisfies $p_{i}=P_{\text{M}}(t_{i}\mid\mathcal{Z},P,t_{<i},t_{<i}\in S)$ and captures how likely the frozen model is to emit token $t_{i}$ when conditioned on the soft prompt and the observed prefix tokens.
@@
Additive loss penalizes the model more for specific error-prone tokens. We evaluated the list of error-prone tokens from training, and we found that the majority of the error-prone tokens are more commonly used tokens, and they are widely generalized in various usages. In that case, we count these tokens and include them in the $K_{\mathcal{M}}$ set as the basis for additive loss.
@@
$K_{\mathcal{M}}$ represents the total number of error-prone tokens, $x_{i}$ denotes the $i_{th}$ input example, and $\mathcal{M}(x_{i})$ is the set of error-prone tokens in $x_{i}$. The term $\log P_{\text{M}}(t\mid\mathcal{Z},P)$ is the log probability of token $t$ given the context $\mathcal{Z}$ and model parameters $P$. The scaling factor $\alpha$ controls the weight of the additive loss, ensuring the penalty is proportionate. By summing over these tokens and averaging over $K_{\mathcal{M}}$, the model receives a consistent and normalized penalty, encouraging it to improve predictions for commonly wrongly predicted tokens, ultimately leading to better performance and generalization.
@@
Based on the analysis of $L_{\text{Addi}}$ and the frequency of token errors, we found that frequently mispredicted tokens are often among the most commonly used tokens. This led to Common loss, which penalizes these frequently used tokens to improve overall performance.
@@
Eq. 4 normalizes the loss by the number of elements in $K_{\mathcal{C}}$, representing the top 10% most frequently used tokens. It sums the weighted log-probabilities of tokens $t$ in the context $\mathcal{C}(x_{i})$ of each element $x_{i}$. Here, $\beta$ is a weighting factor adjusting the influence of each term. $K_{\mathcal{C}}=\{k\in K_{\text{Vocab}}\mid\text{rank}(k)\leq 0.1\times|K_{\text{Vocab}}|\}$ defines $K_{\mathcal{C}}$ as the set of tokens in the vocabulary $K_{\text{Vocab}}$ that are within the top 10% in terms of usage frequency, with $\text{rank}(k)$ providing the frequency rank and $|K_{\text{Vocab}}|$ being all tokens in the model tokenizer’s vocabulary.
@@
Equation 5 computes the absolute deviation between each candidate sequence’s diversity and the desired target. The diversity score for each sequence $\mathcal{D}(S_{i}^{\prime})$ is calculated by first excluding the end-of-sequence tokens, denoted as EOS. For each sequence, we consider the set of unique tokens $\{S_{i,j}\}$, where $S_{i,j}$ denotes the $j_{\text{th}}$ token in the $i_{\text{th}}$ sequence. The diversity of a sequence is then determined by the ratio of the number of unique tokens to the total number of tokens in the sequence, excluding EOS tokens. The diversity score is calculated as the absolute difference between this diversity ratio and the desired diversity value, referred to as $\mathcal{D}_{\text{Optimal}}$. After computing the diversity scores for all sequences, the sequences are sorted based on these scores. The sequences with diversity scores closest to the optimal value are considered the most consistent. By focusing on consistency, Self Consistency reduces the likelihood of incoherent or irrelevant text generation. Also, using multiple candidate sequences and a consistency check ensures the selected sequence is the most reliable.
@@
We utilize two datasets from the LM Extraction Benchmark
(Carlini et al. 2024), originally from The Pile (Gao et al. 2020). These datasets consist of NumPy token arrays, tokenized using the GPT-2 tokenizer, which is compatible with GPT-Neo. Dataset $\mathcal{D}_{1}$ contains 15,000 sequences (50-token prefix and 50-token suffix), used by Ethicist (Zhang, Wen, and Huang 2023); Dataset $\mathcal{D}_{2}$ has 16,000 sequences (150-token prefix and 50-token suffix), partially used by CLM (Ozdayi et al. 2023).
@@
We extend CoSPED to evaluate across different model architectures.
GPT-Neo
Previous works (Ozdayi et al. 2023; Zhang, Wen, and Huang 2023) commonly select GPT-Neo 1.3B, trained on The Pile, as it shares the tokenizer used in LM Extraction Benchmark. Its widespread use makes it a suitable baseline for comparison.
Pythia
We further evaluate CoSPED on the Pythia 1.4B model (Biderman et al. 2023), a model family trained on The Pile with open weights and reproducible settings. We note that prior works have not explored CoSPED-like extraction tasks on Pythia.
To adapt CoSPED, we implemented a padding strategy for input alignment across models with different tokenizer vocab sizes. Our preprocessing ensures compatibility without affecting output validity.
@@
The experiment results presented in Table 1 provide a comparison of various loss components, evaluated using two metrics: $\text{ER}_{50}$ and $\text{ER}_{30}$. Each result represents the average of five repeated sessions. This experiment is conducted using Dataset $\mathcal{D}_{1}$.
The loss components explored in this study vary between base loss functions and loss components, including our proposed losses and existing losses. Our base loss choices are MLE and Focal Loss, both currently active losses. Our loss components are Smooth Loss, Dynamic Loss, Additive Loss, and Common Loss. Smooth loss is a preexisting loss proposed by (Zhang, Wen, and Huang 2023), and the others are our novel proposal.
@@
Table 2: CoSPED Experiment Results of Decoding Strategy
@@
Soft Prompt
Length
Prefix ER50 ER30 ER10
CoSPED Ethicist CLM CoSPED Ethicist CLM CoSPED Ethicist CLM
20 50 55.7 53.2 50.0 62.7 60.1 57.0 74.3 73.3 71.8
100 79.5 77.9 75.8 83.4 81.8 79.9 91.2 90.0 88.6
150 94.8 93.0 91.9 96.3 95.7 94.1 98.3 97.9 97.1
100 50 64.5 61.2 53.6 68.9 66.5 59.7 78.2 76.8 74.3
100 80.8 79.3 76.2 84.3 83.0 80.5 91.5 90.6 89.0
150 95.1 94.5 92.5 97.1 96.4 94.4 98.6 97.9 96.8
150 50 62.1 59.9 54.7 67.0 64.7 60.1 79.0 77.8 73.1
100 80.6 80.2 79.5 85.0 84.7 84.2 92.1 92.0 91.7
150 93.9 93.6 91.0 95.9 95.0 93.1 98.3 98.2 97.3
Decoding Strategy
Beam Search achieves the highest $\text{ER}_{50}$ at 64.5%, but our Self Consistency strategy achieves comparable results ($\text{ER}_{50}$) at 64.2% with better efficiency (13.4ms vs 18.7ms per generation). Self Consistency evaluates consistency across multiple outputs, providing faster generation without compromising quality.
Table 4: Experiment Results of Methods and Baselines

### v3

@@ Abstract
Large language models have gained widespread attention recently, but their potential security vulnerabilities, especially privacy leakage, are also becoming apparent. To test and evaluate for data extraction risks in LLMs, we propose CoSPED, short for Consistent Soft Prompt Targeted Data Extraction and Defense. We introduce several innovative components, including Dynamic Loss, Additive Loss, Common Loss, and Self Consistency Decoding Strategy, and tested to enhance the consistency of the soft prompt tuning process. Through extensive experimentation with various combinations, we achieved an extraction rate of 65.2% at a 50-token prefix comparison. Our comparisons of CoSPED with other reference works confirm our superior extraction rates. We evaluate CoSPED on more scenarios, achieving Pythia model extraction rate of 51.7% and introducing cross-model comparison. Finally, we explore defense through Rank-One Model Editing and achieve a reduction in the extraction rate to 1.6%, which proves that our analysis of extraction mechanisms can directly inform effective mitigation strategies against soft prompt-based attacks.
footnotetext: Accepted by the AAAI 2026 Conference (Special Track on AI Alignment). This is an extended version with appendices.
@@ Introduction
With the growing popularity of ChatGPT (Wu et al. 2023; Yan et al. 2023) and other Large Language Models (LLMs) (Devlin et al. 2019; Radford et al. 2019), their remarkable capabilities come with emerging security and privacy risks (Carlini et al. 2021; Yao et al. 2024). Recent works indicate that LLMs tend to memorize training data, resulting in considerable potential for leakage (Carlini et al. 2022b), which raises an urgent need for deeper analysis and effective mitigation strategies.
@@
Soft prompts (Lester, Al-Rfou, and Constant 2021) offer a light-weight handle for such attacks because they guide model behavior without changing weights. The Controllable Language Model (CLM) framework (Ozdayi et al. 2023) offers an initial treatment of this problem by training soft prompts for extraction attacks. Ethicist (Zhang, Wen, and Huang 2023) strength this area with smoothing loss and calibrated confidence estimation to prioritize high-risk tokens.
From the analysis of previous works, we observe that generation-based extraction methods often produce highly variable or incoherent outputs across repetitions, loss configurations, and decoding choices. These findings reveal three concrete gaps: (1) Most approaches optimize a simple loss in isolation, so gradients are dominated by noisy tokens and different random seeds lead to inconsistent prompts. (2) Decoding strategies are typically selected heuristically and decoupled from the loss design, causing low consistency between generated outputs and memorized training data. (3) Evaluations usually stay within one model family and rarely connect extraction behavior back to deployable defenses, limiting our ability to translate insights into mitigation.
@@
  • We conduct extensive studies on multiple open-source LLMs, including GPT-Neo series and Pythia series, introducing model structural related influence.
  • We implement and evaluate a defense mechanism based on the ROME [^1] ROME: https://rome.baulab.info/ model editing, demonstrating its effectiveness in mitigating soft prompt-based extraction attacks.
@@
Large-scale pre-trained language models often unintentionally memorize training data, leading to privacy risks where sensitive information can be leaked (Carlini et al. 2021, 2022b; Tirumala et al. 2022). Recent surveys and large-scale extraction pipelines show that this vulnerability persists even in production systems (Ishihara 2023; Nasr et al. 2025) and can be amplified through decomposition-based decoding or other attack heuristics (Su et al. 2024; Yu et al. 2023). These works emphasize that extraction robustness must be evaluated beyond a single model or decoding configuration.
Various studies have shown that models can inadvertently store and reveal personal information, which can be extracted through targeted queries (Carlini et al. 2019; Lehman et al. 2021). Defense efforts include data deduplication (Kandpal, Wallace, and Raffel 2022), sensitive information deletion (Patil, Hase, and Bansal 2024), and differential privacy (Abadi et al. 2016; McMahan et al. 2018), but these strategies remain insufficient against evolving attacks (Yao et al. 2024). Thus, frameworks are required that can rigorously evaluate extraction across various scenarios.
Extraction attacks focus on recovering memorized data, including model stealing (Kariyappa, Prakash, and Qureshi 2021; Truong et al. 2021), gradient leakage (Li et al. 2023), and training data extraction (Ozdayi et al. 2023; Yang et al. 2024). Membership inference attacks (Shokri et al. 2017), on the other hand, attempt to determine whether a particular data record was used in the training dataset of a model(Carlini et al. 2022a). While membership inference attacks determine whether a point was used for training (Truex et al. 2019; Carlini et al. 2022a; Fu et al. 2025), training-data extraction aims to reconstruct the underlying content, making CoSPED complementary to but distinct from membership-focused prompt tuning.
@@
Prompt tuning (Lester, Al-Rfou, and Constant 2021) optimizes continuous input embeddings (soft prompts) to guide model behavior without modifying model parameters (Li and Liang 2021). Unlike text prompts, soft prompts directly manipulate input at the embedding level.
Recent works have exploited soft prompts for data extraction attacks. CLM (Ozdayi et al. 2023) trains soft prompts to extract memorized suffixes given prefixes, achieving improved extraction rates over baselines. Ethicist (Zhang, Wen, and Huang 2023) enhances extraction with a smoothing loss that optimizes generation probabilities for high-loss tokens and uses calibrated confidence estimation for suffix selection. Follow-up efforts propose decoding heuristics and prompt-inversion tricks to further stabilize extraction (Yu et al. 2023; Zhang, Carlini, and Ippolito 2024). However, these methods still suffer from high variance across runs, focus on a single model family, or do not offer a unified view that connects extraction and defense.
Overall, prior works have primarily focused on simple loss optimization and heuristic decoding within a single model family. In the following section, we introduce CoSPED, a consistency-driven framework that addresses these limitations in data extraction and defense.
@@
We begin by formalizing the problem setting and notations. Let $X$ denote a continuous token sequence from the pre-training corpus of a large language model (LLM). Each sequence $X$ is divided into a prefix $P$ containing $k_{P}$ tokens and a suffix $S$ containing $k_{S}$ tokens. We represent training corpus as $D=\{X_{i}(P_{i},S_{i})\}$, where each sequence $X_{i}$ is paired with its corresponding prefix-suffix pair. The attacker-accessible split is denoted as $D_{a}=\{X_{i}(P_{i},S_{i})\mid X_{i},P_{i},S_{i}\in D\}.$ This dataset enumerates all prefix-suffix pairs available for soft prompt tuning, ensuring that optimization covers the same statistical distribution as memorized corpus. We assume $D$ contains commonly used public sources such as Books3222Books3: https://github.com/soskek/bookcorpus/issues/27 (Presser 2020), Common Crawl [^3] Common Crawl: https://commoncrawl.org/ (Common Crawl 2023), and The Pile [^4] The Pile: https://pile.eleuther.ai/ (Gao et al. 2020), making this setup realistic for large-scale pre-trained models. The objective is to recover memorized suffixes $S$ given corresponding prefixes $P$ by optimizing soft prompt without modifying model weights.
@@
Our approach employs a trainable soft prompt $\mathcal{Z}\in\mathbb{R}^{K\times d}$, where $K$ is the prompt length and $d$ is the model’s embedding dimension. During training, $\mathcal{Z}$ is concatenated with prefixes $P_{i}$ sampled from $D_{a}$, and only $\mathcal{Z}$ is optimized to minimize the loss $\mathcal{L}$ for generating the corresponding suffix $S_{i}$, while all LLM parameters remain frozen. After training, the optimized soft prompt is evaluated on a disjoint test split $D_{p}$, where prefixes $P_{p}$ are used to generate predicted suffixes $S_{p}$. The outputs are compared against ground-truth suffixes using exact-match and token-level accuracy metrics to measure extraction success. Figure 2 provides an overview of this pipeline, illustrating the soft prompt initialization, optimization process, and evaluation workflow.
@@
We consider a targeted training-data extraction adversary with white-box access to the victim LLM. The attacker can inspect model parameters, gradients, and internal activations, and is permitted to tune continuous soft prompts $\mathcal{Z}$ while keeping model weights frozen. They may query the model repeatedly but cannot alter pre-training data or access user-specific logs. The goal is to reconstruct memorized suffixes $S$ corresponding to given or guessed prefixes $P$. Our defense evaluation adopts the same assumptions to examine how editing internal representations can reduce extraction success rate while preserving normal downstream utility.
@@
By default, an LLM is trained with a maximum likelihood estimation (MLE) loss. Besides MLE, we explore several additional losses, including smooth loss, focal loss, dynamic loss, additive loss, and common loss, some of which consist of the best $\mathcal{L}_{\text{extra}}$ shown in Fig. 2, which can either enhance or replace the original MLE loss.
$S$ represents the suffix sequence the model is trying to generate, and MLE loss is calculated based on each $t_{i}$ token loss, which is influenced by soft prompts $\mathcal{Z}$, input prefix $P$, and previously generated tokens $t_{<i}$.
Ethicist proposed an extra loss function called Smooth loss, which is meant to increase the top $N$ lowest loss values and additionally optimize the generation probabilities for these $N$ tokens. It gives improvements while combining with original $\mathcal{L}_{\text{MLE}}$, but in our method, we are exploring more combinations of different losses, including smooth loss, to test for more potential.
@@
Focal loss (Lin et al. 2017), equation as $\mathcal{L}_{\text{Focal}}=-\frac{1}{|S|}\sum_{i=1}^{|S|}\alpha_{t}(1-p_{i})^{\gamma}\log p_{i},$ is a modification of standard cross-entropy loss, primarily designed to address the class imbalance problem. In our scenario, focal loss focuses on penalizing hard-to-classify tokens more than easier tokens, thereby emphasizing learning from difficult instances. Focal loss reweights each token-level log likelihood with $\alpha_{t}$ and the focusing parameter $\gamma$ to emphasize hard examples. The probability term satisfies $p_{i}=P_{\text{M}}(t_{i}\mid\mathcal{Z},P,t_{<i},t_{<i}\in S)$ and captures how likely the frozen model is to emit token $t_{i}$ when conditioned on the soft prompt and the observed prefix tokens.
@@
Additive loss penalizes the model more for specific error-prone tokens. We evaluated the list of error-prone tokens from training, and we found the majority of error-prone tokens are more commonly used tokens, and they are widely generalized in various usages. We count these tokens and include them in the $K_{\mathcal{M}}$ set as the basis for additive loss.
@@
$K_{\mathcal{M}}$ represents the total number of error-prone tokens, $x_{i}$ denotes the $i_{th}$ input example, and $\mathcal{M}(x_{i})$ is the set of error-prone tokens in $x_{i}$. The term $\log P_{\text{M}}(t\mid\mathcal{Z},P)$ is the log probability of token $t$ given the context $\mathcal{Z}$ and model parameters $P$. Scaling factor $\alpha$ controls the weight of additive loss, ensuring the penalty is proportionate. By summing over these tokens and averaging over $K_{\mathcal{M}}$, the model receives a consistent and normalized penalty, encouraging it to improve predictions for commonly wrongly predicted tokens, leading to better performance and generalization.
@@
Based on the analysis of $L_{\text{Addi}}$ and the frequency of token errors, we found frequently mispredicted tokens are often among the most commonly used tokens. This led to Common loss, penalizing frequently used tokens to improve overall performance.
@@
Eq. 3 normalizes the loss by the number of elements in $K_{\mathcal{C}}$, representing the top 10% most frequently used tokens. It sums the weighted log-probabilities of tokens $t$ in the context $\mathcal{C}(x_{i})$ of each element $x_{i}$. Here, $\beta$ is a weighting factor adjusting the influence of each term. $K_{\mathcal{C}}=\{k\in K_{\text{Vocab}}\mid\text{rank}(k)\leq 0.1\times|K_{\text{Vocab}}|\}$ defines $K_{\mathcal{C}}$ as the set of tokens in the vocabulary $K_{\text{Vocab}}$ that are within top 10% in terms of usage frequency, with $\text{rank}(k)$ providing frequency ranks and $|K_{\text{Vocab}}|$ being all tokens in model tokenizer’s vocabulary.
@@
Equation 4 computes the absolute deviation between each candidate sequence’s diversity and the desired target. The diversity score for each sequence $\mathcal{D}(S_{i}^{\prime})$ is calculated by first excluding the end-of-sequence tokens, denoted as EOS. For each sequence, we consider the set of unique tokens $\{S_{i,j}\}$, where $S_{i,j}$ denotes the $j_{\text{th}}$ token in the $i_{\text{th}}$ sequence. The diversity of a sequence is then determined by the ratio of the number of unique tokens to the total number of tokens in the sequence, excluding EOS tokens. The diversity score is calculated as the absolute difference between this diversity ratio and the desired diversity value, referred to as $\mathcal{D}_{\text{Optimal}}$. After computing the diversity scores for all sequences, the sequences are sorted based on these scores. The sequences with diversity scores closest to the optimal value are considered the most consistent. By focusing on consistency, Self Consistency reduces the likelihood of incoherent or irrelevant text generation. Also, using multiple candidate sequences and a consistency check ensures that the selected sequence is the most reliable.
@@
We utilize two datasets from the LM Extraction Benchmark [^5] LM Extraction Benchmark: https://github.com/google-research/lm-extraction-benchmark (Carlini et al. 2024), originally from The Pile (Gao et al. 2020). These datasets consist of NumPy token arrays, tokenized using the GPT-2 tokenizer, which is compatible with GPT-Neo. Dataset $\mathcal{D}_{1}$ contains 15,000 sequences (50-token prefix and 50-token suffix), used by Ethicist (Zhang, Wen, and Huang 2023); Dataset $\mathcal{D}_{2}$ has 16,000 sequences (150-token prefix and 50-token suffix), partially used by CLM (Ozdayi et al. 2023).
@@
We extend CoSPED to evaluate across different models. Previous works (Ozdayi et al. 2023; Zhang, Wen, and Huang 2023) commonly select GPT-Neo 1.3B [^6] GPT-Neo 1.3B: https://huggingface.co/EleutherAI/gpt-neo-1.3B, trained on The Pile, as it shares tokenizer used in LM Extraction Benchmark. Its widespread usage makes it fit for baseline comparison.
We further evaluate CoSPED on Pythia 1.4B model [^7] Pythia 1.4B: https://huggingface.co/EleutherAI/pythia-1.4b (Biderman et al. 2023). Pythia is a model family trained on The Pile with open weights and reproducible settings. We note that prior works have not explored CoSPED-like extraction tasks on Pythia. To adapt CoSPED, we implemented a padding strategy for input alignment across models with different tokenizer vocab sizes. Our preprocessing ensures compatibility without affecting output validity.
@@
The experiment results presented in Table 1 provide a comparison of various loss components, evaluated using two metrics: $\text{ER}_{50}$ and $\text{ER}_{30}$. Each result represents the average of five repeated sessions. This experiment is conducted using Dataset $\mathcal{D}_{1}$. The loss components explored in this study vary between base loss functions and loss components, including our proposed losses and existing losses. Base loss choices are MLE and Focal Loss, both commonly used losses. Loss components are Smooth Loss, Dynamic Loss, Additive Loss, and Common Loss. Smooth loss is a preexisting loss by Ethicist, and the others are our novel proposal.
@@
These effects align with our consistency-oriented objective. Smooth loss maintains pressure on a fixed set of hard tokens, whereas Dynamic loss expands the set whenever the current batch becomes harder, preventing variance spikes between runs. Additive loss denoises the gradient by emphasizing tokens that repeatedly cause mistakes, so the dynamic schedule does not chase incidental noise. Finally, combining Focal loss with mild label smoothing sharpens attention on truly difficult tokens while avoiding overconfidence on easy ones, which stabilizes optimization and sets up the self-consistency decoding with less noise.
@@
Decoding Strategy
Beam Search achieves the highest $\text{ER}_{50}$ at 64.5%, but our Self Consistency strategy achieves comparable results on $\text{ER}_{50}$ at 64.2% with better efficiency (13.4ms vs 18.7ms per generation). Self Consistency evaluates consistency across multiple outputs, providing faster generation without compromising quality.
Soft Prompt Length Prefix ER50 ER30 ER10
CoSPED Ethicist CLM CoSPED Ethicist CLM CoSPED Ethicist CLM
20 50 55.7 53.2 50.0 62.7 60.1 57.0 74.3 73.3 71.8
100 79.5 77.9 75.8 83.4 81.8 79.9 91.2 90.0 88.6
150 94.8 93.0 91.9 96.3 95.7 94.1 98.3 97.9 97.1
100 50 64.5 61.2 53.6 68.9 66.5 59.7 78.2 76.8 74.3
100 80.8 79.3 76.2 84.3 83.0 80.5 91.5 90.6 89.0
150 95.1 94.5 92.5 97.1 96.4 94.4 98.6 97.9 96.8
150 50 62.1 59.9 54.7 67.0 64.7 60.1 79.0 77.8 73.1
100 80.6 80.2 79.5 85.0 84.7 84.2 92.1 92.0 91.7
150 93.9 93.6 91.0 95.9 95.0 93.1 98.3 98.2 97.3
@@
Table 3: CoSPED Experiment Results of Decoding Strategy
Comparison against Prior Works
Discussion
CoSPED VS. Ethicist and CLM